# Detection of Intelligent Tampering in Wireless Electrocardiogram Signals Using Hybrid Machine Learning

### v1

@@ VII Conclusions
This research evaluated the performance of CNN, ResNet, and hybrid deep learning architectures for ECG tampering detection, alongside a Siamese Transformer-based model for person-specific ECG verification. Experimental results show that in highly fragmented manipulation scenarios, CNN, FeatCNN-TranCNN, FeatCNN-Tran and ResNet models achieved above 99.5% accuracy; for subtle manipulations (e.g., 50%-50%) and (75%-25%) replacement, the FeatCNN-TranCNN models performed consistently better, achieving above 98% accuracy. The performance of CNN-based models (standard CNN and ResNet) is high in moderately tampered conditions; however, they exhibited high variability during training, indicating sensitivity to hyperparameters and a tendency toward overfitting. In contrast, the hybrid model (transformer-CNN) achieved high performance and also demonstrated more stable training across varying tampering levels. Pure Transformer models (Tran-DeepFFN) that work on CWT-transformed inputs performed relatively less well in terms of accuracies, particularly in cases of subtle tampering scenarios. Additionally, CWT-FeatCNN-Tran; where the inputs are first modified by CWT achieved moderate accuracies (about 95%-96%). It is also worth mentioning that CWT transformations also added additional computational overhead at the pre-processing step, and higher-dimensional CWT inputs (2048×96) increased model complexity and inference time. Hybrid designs that incorporated CNN components for feature extraction or into the Transformer feed-forward network performed better in terms of finding a balance between stability, computational efficiency, and peak accuracy. For person-specific ECG verification, the pure Transformer-Siamese network achieved an accuracy of 98.30%. In contrast, the hybrid CNN$+$Transformer-Siamese model delivered perfect verification performance with 100% accuracy. Future work will focus on incorporating more dataset to include varied patient profiles and integrating the model developed with edge computing platforms to enable real-time detection in wearable and mobile devices.

### v2

@@ VII Conclusions
This research evaluated the performance of CNN, ResNet, and transformer-based hybrid models for ECG tampering detection. Experimental results show that in highly fragmented manipulation scenarios, CNN, FeatCNN-TranCNN, FeatCNN-Tran and ResNet models achieved above 99.5% accuracy; for subtle manipulations (e.g., 50%-50%) and (75%-25% replacement), the FeatCNN-TranCNN models performed consistently better, achieving above 98% accuracy. The performance of CNN-based models (standard CNN and ResNet) is high in moderately tampered conditions; however, they exhibited high variability during training, indicating sensitivity to hyperparameters and a tendency toward overfitting. In contrast, the hybrid model achieved high performance and also demonstrated more stable training across varying tampering levels. Pure Transformer models (Tran-DeepFFN) that work on CWT-transformed inputs performed relatively poorly in terms of accuracy, particularly in cases of subtle tampering scenarios. Additionally, CWT-FeatCNN-Tran; where the inputs are first modified by CWT, achieved moderate accuracies (about 95%-96%). It is also worth mentioning that CWT transformations also added additional computational overhead at the pre-processing step, and higher-dimensional CWT inputs (2048×96) increased model complexity and inference time. Hybrid designs that incorporated CNN components for feature extraction or into the Transformer feed-forward network performed better in terms of finding balance between stability, computational efficiency, and peak accuracy. Future work will focus on incorporating more dataset to include varied patient profiles and integrating the models with edge computing platforms to enable real-time detection in wearable and mobile devices.